OpenStudy (anonymous):

one countrys GDP is $100 in year 0. GDP grows every year by 2.4. what is next 20 years GDP

OpenStudy (anonymous):

y=2.4x+100 y=2.4(20)+100 y=48+100 y=148

OpenStudy (anonymous):

but i thought i need to use compound interest equation?